1.Draw the pH titration curve for the titration of 35 mL of 0.150 M acetic acid with 0.200 m Include the pH values and NaOH vI'm not sure how to solve this! Any help is appreciated!  
Answer #1

1) the initial pH

Initially there is only acetic acid present in the solution, hence pH of solution will be due to dissociation of acetic acid. Consider a dissociation of acetic acid in water.

CH3COOH (aq) + H2O (l)phpDQGRLY.png CH3COO -(aq) + H3O +(aq)  

For above reaction, dissociation constant is Ka = [CH3COO - ] [H3O + ] / [CH3COOH] = 1.8 \times 10 -05

Consider X moles of acetic acid dissociated at equilibrium, then we can write

K a = (X) \times ( X) / 0.150 - X = 1.8 \times 10 -05

Acetic acid is a weak acid. Hence, we can assume X is very small as compared to 0.150. Hence we can write 0.150 - X phpoF7bep.png 0.150.

X 2 / 0.150 = 1.8 \times 10 -05

  X 2 = 0.150 ( 1.8 \times 10 -05 )

  X 2 = 2.7 \times 10 -06

X = 1.64 \times 10 -03 M = [CH3COO - ] =[H3O + ]

We have, pH = -log [H3O + ] = - log 1.64 \times 10 -03 = 2.78

2) pH of solution after addition of 15.0 ml NaOH

Consider a reaction , CH3COOH (aq) + NaOH (aq)phpt5MS4K.png  CH3COONa (aq) + H2O (l)

mmol of CH3COOH = concentration x volume = 0.150 \times 35 =5.25 mmol

mmol of NaOH = 0.200 \times 15.0 = 3.00 mmol

Excess mmol of CH3COOH = initial mmol - mmol after reaction with NaOH = 5.25 - 3.00=2.25

mmol of CH3COONa produced = mmol of NaOH added to solution = 3.00

Volume of solution =35.0 ml +15.0 ml = 50.0 ml

[CH3COOH]= No of moles / Volume of solution in L

[CH3COOH] = [2.25 /1000] / [50.0 /1000]

[CH3COOH ] = 0.045 M

[CH3COONa]= [3.00 /1000] / [50.0 /1000]

[CH3COONa]= 0.0600 M

At this stage, solution contain weak acid (acetic acid) and its salt (sodium acetate) .Hence, solution acts as a buffer solution and it's pH is calculated by using Henderson's equation

pH = pKa + log [ salt] / [ Acid]

pH = - log Ka + log [CH3COONa] / [CH3COOH]

pH = -log (1.8 x 10 -05) + log 0.0600 /0.045

pH = 4.74 + 0.125

pH = 4.86

3) the volume of NaOH at the equivalence point

Consider a reaction , CH3COOH (aq) + NaOH (aq)phpt5MS4K.png  CH3COONa (aq) + H2O (l)

From reaction, 1 mol CH3COOH \equiv 1 mol NaOH

We have relation, M acidphpw283bS.png V acid = M basephptMBdZL.png V base

   V base = M acidphpGCf8ms.png V acid / M base

   V base =0.150 M phpm0P3EE.png 35.0 ml / 0.200 M

   V base = 26.25 ml

Volume of NaOH at equivalence point = 26.25 ml

4) the pH at the equivalence point

At equivalence point, all acetic acid is consumed by added NaOH. pH of solution will be due to dissociation of CH3COONa in water.

CH3COO- (aq) + H2O (l)phpV8JN8M.png CH3COOH (aq) + OH -(aq)

Kb = [CH3COOH] [OH - ] / [CH3COO- ]= Kw / Ka = 10 -14 / 1.8 \times10 -05 = 5.56 \times 10 -10

mmol of CH3COONa produced = mmol of NaOH added to the solution =0.200 \times 26.25  =5.25 mmol

Volume of the solution = volume of acetic acid + volume of NaOH = 35.0 ml +26.25 ml  = 61.25 ml

[CH3COONa]= [5.25/1000] / [61.25 /1000] = 0.0857 M

Consider X moles of sodium acetate dissociated at equilibrium, then we can write

Kb = (X ) ( X) / 0.0857 - X = 5.56 \times 10 -10

Here X is very small as compared to 0.857. Hence we can write 0.0857 - X php3jmClp.png 0.0857.

X 2 = 0.0857 ( 5.56 \times 10 -10 )  

X 2 = 4.76\times 10 -11

X = 6.90 \times10 -06 M = [ OH - ]

We know that , pOH = -log [ OH - ] = - log 6.90 \times10 -06 = 5.16

We have relation, pH + pOH = 14

Therefore, pH = 14 - pOH = 14 - 5.16 = 8.84

5) pH after 50.0 ml NaOH added

After equivalence point, pH of solution will be due to NaOH.

mmol of NaOH =0.200 \times 50.0 = 10.0 mmol

mmol of CH3COOH =0.150 \times 35.0 = 5.25 mmol

mmol of excess NaOH = 10.0 - 5.25 = 4.75 mmol

Volume of solution = volume of acid + volume of base = 35.0 + 50.0 = 85.0 ml

[NaOH ] = [ 4.75 /1000] / [ 85.0 /1000] = 0.05588 M

We know that , pOH = -log [ OH - ] = - log 0.0558 = 1.31

We have relation, pH + pOH = 14

Therefore, pH = 14 - pOH = 14 - 1.31 = 12.69

due to Na on pH due to CH3C001 20 2K 2 Sk 35 40 45 46 se NON NON- NAOH GE​​​​​​​